If there were 60,000 pounds of raw materials on hand on January 1, 120,000 pounds are...

  1. If there were 60,000 pounds of raw materials on hand on January 1, 120,000 pounds are desired for inventory at January 31, and 360,000 pounds are required for January production, how many pounds of raw materials should be purchased in January?

    A

    300,000 pounds

    B

    480,000 pounds

    C

    240,000 pounds

    D

    420,000 pounds

Strand Company is planning to sell 200 buckets and produce 190 buckets during March. Each bucket requires 500 grams of plastic and one-half hour of direct labor. Plastic costs $10 per 500 grams and employees of the company are paid $15.00 per hour. Manufacturing overhead is applied at a rate of 110% of direct labor costs. Strand has 300 kilos of plastic in beginning inventory and wants to have 200 kilos in ending inventory. How much is the total amount of budgeted direct labor for March?

A

$1,500

B

$3,000

C

$1,425

D

$2,850

Teller Co. is planning to sell 600 boxes of ceramic tile, with production estimated at 580 boxes during May. Each box of tile requires 44 pounds of clay mix and a quarter hour of direct labor. Clay mix costs $0.40 per pound and employees of the company are paid $12.00 per hour. Manufacturing overhead is applied at a rate of 110% of direct labor costs. Teller has 2,600 pounds of clay mix in beginning inventory and wants to have 3,000 pounds in ending inventory.

What is the total amount to be budgeted for manufacturing overhead for the month?

A

$1,914

B

$1,980

C

$7,656

D

$7,920
Answer #1

Answer:

Option (D) 420,000 pounds

Pounds of materials to be purchased

Consumption = Opening + Purchasing - Closing
360000= 60000+ Purchases - 120000
Purchases = 360000-60000+120000
Purchases = 420,000

Hence pounds of materials to be purchased in the month of January = 420,000

190 buckets*1/2 hour per bucket*$15.00 per hour = 1425
Answer: Option (C) $1425

Direct labor cost = 580 boxes *.25 Hours * 12 = $ 1740
Manufacturing overhead Budgeted = 1740 *110% = $ 1,914
Answer:
Option (A) $1,914